OpenStudy (qwerd722):

Integral Calculus (Volumes of Revolution) Find the volume generated when the area bounded by the curve y=x^4, the y-axis, and y=16. is revolved about the y-axis in the first quadrant.

OpenStudy (qwerd722):

Graph

OpenStudy (kikuo):

What grade is this?

OpenStudy (mww):

well you can use either slices or cylindrical shells. Slices might be easier for this case. Best you draw up the question first like so:  So we see each slice/disc area is given by \[A = \pi r^2 = \pi x^2\] As the depth of each slice/disc is change in y units We have \[V = \lim_{\Delta y \rightarrow 0}\sum_{0}^{16} \pi x^2 \Delta y = \int\limits_{0}^{16}\pi x^2 dy\] The latter part of this equation you should be familiar with already Sub in x into this formula (in terms of y and you should evaluate the volume of revolution.